### v2

@@ 5 Conclusions
In this paper, we study optimal stopping for diffusion processes in the exploratory RL framework of Wang et al., (2020). Different from Dong, (2024) and Dai and Dong, (2024) which directly randomize stopping times for exploration through mixed strategies characterized by Cox processes, we turn the problem into stochastic control by approximating the variational inequality problem with the so-called penalized PDE. This enables us to apply all the available results and methods developed recently in RL for controlled diffusion processes (Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou,, 2023). Our simulation study shows that the resulting RL algorithms achieve comparable, and sometimes better, performance in terms of accuracy and speed than those in Dong, (2024) and Dai and Dong, (2024).
@@
There are many interesting problems to investigate along the lines of this paper, including convergence and regret analysis, choice of the temperature parameter, and more complex underlying processes to stop.
@@ Appendix A Proofs of statements
Suppose there is a positive sequence $\{K_{n}\}_{n\geqslant 1}$ that increases to infinity, and $v^{(n)}$ solves (2.5) with $K=K_{n}$. Denote by $a_{n}=v_{t}^{(n)}+\mathscr{L}v^{(n)},\ b_{n}=g-v^{(n)}$, $a_{\infty}=v_{t}+\mathscr{L}v$, and $b_{\infty}=g-v$, where $v$ is the limit of $v^{(n)}$. From (2.5) we have $a_{n}=-K_{n}(b_{n})^{+}\leqslant 0$. Because $\left(a_{n},b_{n}\right)$ converges to $\left(a_{\infty},b_{\infty}\right)$ as $n\rightarrow\infty$, we get $a_{\infty}\leqslant 0$. On the other hand, on any compact region, $\{a_{n}\}_{n\geqslant 1}$ is locally bounded; so $(b_{n})^{+}=-\frac{a_{n}}{K_{n}}\rightarrow 0,\text{as $n\rightarrow\infty$}$, yielding $b_{\infty}\leqslant 0$.
@@
For notational simplicity, we only consider the case when $X$ and $W$ are both one-dimensional.
@@
This means that $v$ is an upper bound of the value function of problem (2.7). All the inequalities above become equalities if we take $u\in U$ that solves the optimization problem in (2.9). Therefore, we conclude $v$ is the value function of (2.7).
@@
where the last inequality is from the definition of $\tilde{\pi}$. Besides, $\Delta(T,x)=0$ at the terminal time $T$. Therefore, 0 is a sub-solution of $-\partial_{t}u(t,x)-\mathscr{L}u(t,x)+K\tilde{\pi}(t,x)u(t,x)+H\left(t,x,\pi,J^{\pi}\right)-H\left(t,x,\tilde{\pi},J^{\pi}\right)=0$. By the maximum principle, we have $\Delta(t,x)\geqslant 0$, resulting in $J^{\tilde{\pi}}(t,x)\geqslant J^{\pi}(t,x)$.
@@ Appendix D RL for Dynkin games
$-w_{t}-\mathcal{L}w-K\left(w-L(x)\right)^{-}+K\left(w-U(x)\right)^{+}=0,$ (D.3)
@@ Appendix E Finite difference method
$-\mathcal{L}_{BS}u(t,y)=K\left(g(e^{y})-u(t,y)\right)^{+}.$ (E.4)
@@
$-\frac{u_{i+1,j}-u_{i,j}^{k+1}}{\Delta t}-\beta u_{i,j+1}^{k+1}+\left(\alpha+\beta+\rho\right)u_{i,j}^{k+1}-\alpha u_{i,j-1}^{k+1}=K\left(g(e^{y_{j}})-u_{i,j}^{k+1}\right)1_{g(e^{y_{j}})-u_{i,j}^{k}>0},$
@@
$-\beta u_{i,j+1}^{k+1}+\left(\frac{1}{\Delta t}+\alpha+\beta+\rho+K1_{g(e^{y_{j}})-u_{i,j}^{k}>0}\right)u_{i,j}^{k+1}-\alpha u_{i,j-1}^{k+1}=\frac{u_{i+1,j}}{\Delta t}+Kg(e^{y_{j}})1_{g(e^{y_{j}})-u_{i,j}^{k}>0}.$ (E.5)
@@
             if ${||u_{i}^{k+1}-u_{i}^{k}||_{\infty}}<\varepsilon*{\max\left\{1,\ ||u_{i}^{k}||_{\infty}\right\}}$ then

### v3

@@ 6 Conclusions
In this paper, we study optimal stopping for diffusion processes in the exploratory RL framework of Wang et al., (2020). Different from Dong, (2024) and Dai and Dong, (2024) which directly randomize stopping times for exploration through mixed strategies characterized by Cox processes, we turn the problem into stochastic control by approximating the variational inequality with the penalized PDE. This enables us to apply all the available results and methods developed recently in RL for controlled diffusion processes (Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou,, 2023). Our simulation study shows that the resulting RL algorithms achieve comparable, and sometimes better, performance in terms of accuracy and speed than those in Dong, (2024) and Dai and Dong, (2024). More importantly, we show experimentally that the proposed algorithms also work well for high-dimensional problems.
There is a notable limitation of the current approach especially to high-dimensional problems. In our implementation, we use neural networks to approximate value functions, which enables us to overcome the curse of dimensionality. However, a challenge may arise if the value functions—such as discontinuous ones—cannot be well approximated by neural networks. Overcoming the difficulty is an interesting yet important question.
@@
There are many other interesting problems to investigate along the lines of this paper, including a more substantial convergence and regret analysis, choice of the temperature parameter, and more complex underlying processes to stop.
@@ Appendix A Proofs of statements
Suppose there is a positive sequence $\{K_{n}\}_{n\geqslant 1}$ that increases to infinity, and $v^{(n)}$ solves (2.5) with $K=K_{n}$. Denote $a_{n}=v_{t}^{(n)}+\mathscr{L}v^{(n)},\ b_{n}=g-v^{(n)}$, $a_{\infty}=v_{t}+\mathscr{L}v$, and $b_{\infty}=g-v$, where $v$ is the limit of $v^{(n)}$. From (2.5) we have $a_{n}=-K_{n}(b_{n})^{+}\leqslant 0$. Because $\left(a_{n},b_{n}\right)$ converges to $\left(a_{\infty},b_{\infty}\right)$ as $n\rightarrow\infty$, we get $a_{\infty}\leqslant 0$. On the other hand, on any compact region, $\{a_{n}\}_{n\geqslant 1}$ is locally bounded; so $(b_{n})^{+}=-\frac{a_{n}}{K_{n}}\rightarrow 0,\text{as $n\rightarrow\infty$}$, yielding $b_{\infty}\leqslant 0$.
@@
For notational simplicity, without loss of generality we only consider the case when $X$ and $W$ are both one-dimensional.
@@
This means that $v$ is an upper bound of the optimal value function of problem (2.8). All the inequalities above become equalities if we take $u\in U$ that solves the optimization problem in (2.10). Therefore, we conclude $v$ is the optimal value function of (2.8).
@@
where the last inequality is from the definition of $\tilde{\pi}$. Besides, $\Delta(T,x)=0$ at the terminal time $T$. Therefore, 0 is a sub-solution of $\partial_{t}u(t,x)-\mathscr{L}u(t,x)+K\tilde{\pi}(t,x)u(t,x)+H\left(t,x,\pi,J^{\pi}\right)-H\left(t,x,\tilde{\pi},J^{\pi}\right)=0$. By the maximum principle, we have $\Delta(t,x)\geqslant 0$, resulting in $J^{\tilde{\pi}}(t,x)\geqslant J^{\pi}(t,x)$.
A.4 Proof of Theorem 3.2
Recall that $v^{\bm{\Pi}}$ is the optimal value function of the exploratory control problem (2.20) with the optimal strategy (2.24). By the same procedure as in the proof of Theorem 3.1, we can show that $v^{\bm{\Pi}}\geqslant\cdots\geqslant J^{\pi^{n+1}}\geqslant J^{\pi^{n}}\geqslant\cdots\geqslant J^{\pi^{0}}$. Next, we show the uniform convergence of the sequence $\left\{J^{\pi^{n}}\right\}_{n=1}^{\infty}$. To proceed, write $\left\|g\right\|_{\infty}=C$ for some positive constant $C$ due to the assumption that $g$ is uniformly bounded.
For any $n\geqslant 1$, denote
$f^{n}(t):=$ $\;\sup_{x}\left\{v^{\bm{\Pi}}(t,x)-J^{\pi^{n}}(t,x)\right\}\geqslant 0,$
$\phi^{n}(t):=$  $\;2K\int_{t}^{T}f^{n}(s)ds,$
$W(t,x):=$ $\;v^{\bm{\Pi}}(t,x)-J^{\pi^{n+1}}(t,x)-\phi^{n}(t).$
Using $\partial_{t}J^{\pi^{n+1}}+\mathscr{L}J^{\pi^{n+1}}+H\left(t,x,\pi^{n+1},J^{\pi^{n+1}}\right)=0$ and $\partial_{t}v^{\bm{\Pi}}+\mathscr{L}v^{\bm{\Pi}}+H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)=0$, we have
$\partial_{t}W+\mathscr{L}W$ $=\partial_{t}v^{\bm{\Pi}}+\mathscr{L}v^{\bm{\Pi}}-\left(\partial_{t}J^{\pi^{n+1}}+\mathscr{L}J^{\pi^{n+1}}\right)-\partial_{t}\phi^{n}$
$=H\left(t,x,\pi^{n+1},J^{\pi^{n+1}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)-\partial_{t}\phi^{n}$
$=H\left(t,x,\pi^{n+1},J^{\pi^{n+1}}\right)-H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)$
$\quad+H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)-\partial_{t}\phi^{n}.$
Noticing that
$H\left(t,x,\pi,J^{\pi}\right)$ $=K\pi(t,x)\left(g(t,x)-J^{\pi}(t,x)\right)-\lambda\mathscr{H}(\pi(t,x)),$
$\mathscr{H}(\pi^{n+1}(t,x))$ $=\pi^{n+1}\log\pi^{n+1}+(1-\pi^{n+1})\log\left(1-\pi^{n+1}\right)$
$=\lambda^{-1}K\left(J^{\pi^{n}}-g\right)\left(1-\pi^{n+1}\right)-\log\left(1+\exp\left(\lambda^{-1}K\left(J^{\pi^{n}}-g\right)\right)\right),$
and
$\mathscr{H}(\pi^{*}(t,x))$ $=\pi^{*}\log\pi^{*}+(1-\pi^{*})\log\left(1-\pi^{*}\right)$
$=\lambda^{-1}K\left(v^{\bm{\Pi}}-g\right)\left(1-\pi^{*}\right)-\log\left(1+\exp\left(\lambda^{-1}K\left(v^{\bm{\Pi}}-g\right)\right)\right),$
we obtain
$H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)$ $=K\pi^{n+1}\left(g-v^{\bm{\Pi}}\right)-K\left(J^{\pi^{n}}-g\right)\left(1-\pi^{n+1}\right)$
$\quad+\lambda\log\left(1+\exp\left(\lambda^{-1}K\left(J^{\pi^{n}}-g\right)\right)\right)$
and
$H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)$ $=K\pi^{*}\left(g-v^{\bm{\Pi}}\right)-K\left(v^{\bm{\Pi}}-g\right)\left(1-\pi^{*}\right)$
$\quad+\lambda\log\left(1+\exp\left(\lambda^{-1}K\left(v^{\bm{\Pi}}-g\right)\right)\right)$
$=K\left(g-v^{\bm{\Pi}}\right)+\lambda\log\left(1+\exp\left(\lambda^{-1}K\left(v^{\bm{\Pi}}-g\right)\right)\right).$
Straightforward calculations show that
$H\left(t,x,\pi^{n+1},J^{\pi^{n+1}}\right)-H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)=K\pi^{n+1}\left(v^{\bm{\Pi}}-J^{\pi^{n+1}}\right)\geqslant 0$
and
$H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)$
$=$ $\ K\left(v^{\bm{\Pi}}-J^{\pi^{n}}\right)\frac{1}{1+\exp\left(-\lambda^{-1}K\left(J^{\pi^{n}}-g\right)\right)}-\lambda\log\left(\frac{1+\exp\left(\lambda^{-1}K\left(v^{\bm{\Pi}}-g\right)\right)}{1+\exp\left(\lambda^{-1}K\left(J^{\pi^{n}}-g\right)\right)}\right).$
It follows from the general inequality $\left|\log(1+\exp(x))-\log(1+\exp(y))\right|\leqslant\left|x-y\right|$ that
$\left|H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)\right|\leqslant 2K\left(v^{\bm{\Pi}}-J^{\pi^{n}}\right)\leqslant 2Kf^{n}(t).$
Thus,
$H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)-\partial_{t}\phi^{n}$
$=$ $\ H\left(t,x,\pi^{n+1},v^{\bm{\Pi}}\right)-H\left(t,x,\pi^{*},v^{\bm{\Pi}}\right)+2Kf^{n}(t)\geqslant 0.$
Consequently,
$\partial_{t}W+\mathscr{L}W\geqslant 0,$
which together with $W(T,x)=0$ implies $W\leqslant 0$ by the maximum principle. Thus, we conclude
$v^{\bm{\Pi}}(t,x)-J^{\pi^{n+1}}(t,x)\leqslant\phi^{n}(t)=2K\int_{t}^{T}f^{n}(s)ds.$
The definition of $f^{n}$ yields
$f^{n+1}(t)\leqslant 2K\int_{t}^{T}f^{n}(s)ds.$
By induction, we have
$f^{n+1}(t)\leqslant\left(2K\right)^{n}\frac{(T-t)^{n}}{n!}\sup_{t}f^{1}(t)\leqslant\frac{(2KT)^{n}}{n!}\sup_{t}f^{1}(t).$
Recall
$J^{\pi}(s,x)=\operatorname{\mathbb{E}}\!\left[\int_{s}^{T}e^{-K\int_{s}^{t}\pi_{v}\operatorname{d}\!v}\big{[}Kg(t,X_{t})\pi_{t}-\lambda\mathscr{H}(\pi_{t})\big{]}\operatorname{d}\!t+e^{-K\int_{s}^{T}\pi_{v}\operatorname{d}\!v}g(T,X_{T})\;\bigg{|}\;X_{s}=x\right],$
$\left\|g\right\|_{\infty}=C$, $e^{-K\int_{s}^{t}\pi_{v}\operatorname{d}\!v}\leqslant 1$ for any $0\leqslant t\leqslant T$, and $0\leqslant\pi_{t}\leqslant 1$ and
$|\mathscr{H}(\pi_{t})|\leqslant\sup_{x\in(0,1)}|x\log x+(1-x)\log(1-x)|=\log 2.$
We thus have
$|J^{\pi}(s,x)|\leqslant\operatorname{\mathbb{E}}\!\left[\int_{s}^{T}\big{[}KC+\lambda\log 2\big{]}\operatorname{d}\!t+C\;\bigg{|}\;X_{s}=x\right]\leqslant(KC+\lambda\log 2)T+C$
for any admissible $\pi$. It hence follows
$\sup_{t}f^{1}(t)=\sup_{t,x}\left\{v^{\bm{\Pi}}(t,x)-J^{\pi^{1}}(t,x)\right\}$ $\leqslant 2((KC+\lambda\log 2)T+C).$
Consequently,
$\left\|v^{\bm{\Pi}}-J^{\pi^{n+1}}\right\|_{\infty}\leqslant\frac{(2KT)^{n}}{n!}2((KC+\lambda\log 2)T+C),$
which converges to 0 as $n$ increases to infinity when $K$ and $\lambda$ are fixed. This concludes the proof.
Finally, let us complement the proof by numerically examining the theoretical convergence rate. As shown in Figure 8, the NN approximation for the stopping value of $(t_{0},x_{0})=(0,40)$ converges to its theoretical value 5.333, which is calculated using the finite difference method, and the relative error decreases exponentially. In implementation, we use stochastic gradient descent to update the NN parameters at each training step, which prevents us from performing the policy evaluation in (3.8) with a complete accuracy. Consequently, in the experiment we do not achieve the $n!$ rate of decay in relative error as predicted by Theorem 3.2. Nevertheless, exponential decay is sufficient to ensure rapid convergence.
Figure 8: Numerical diagnostics for the convergence of the RL algorithms. We use the offline ML Algorithm 1 to learn the stopping value of $(t_{0},x_{0})=(0,40)$, approximated by an NN, for the American put example in Section 4 with $\lambda=0.01$ and $K=50$. The left panel displays the learning curve of the value and the right panel shows that of the relative error, both with respect to training steps. The y-axis of the right panel is plotted on a logarithmic scale.
@@ Appendix D RL for Dynkin games
$w_{t}-\mathcal{L}w-K\left(w-L(x)\right)^{-}+K\left(w-U(x)\right)^{+}=0,$ (D.3)
@@ Appendix E Finite difference method
$\mathcal{L}_{BS}u(t,y)=K\left(g(e^{y})-u(t,y)\right)^{+}.$ (E.4)
@@
$\frac{u_{i+1,j}-u_{i,j}^{k+1}}{\Delta t}-\beta u_{i,j+1}^{k+1}+\left(\alpha+\beta+\rho\right)u_{i,j}^{k+1}-\alpha u_{i,j-1}^{k+1}=K\left(g(e^{y_{j}})-u_{i,j}^{k+1}\right)1_{g(e^{y_{j}})-u_{i,j}^{k}>0},$
@@
$\beta u_{i,j+1}^{k+1}+\left(\frac{1}{\Delta t}+\alpha+\beta+\rho+K1_{g(e^{y_{j}})-u_{i,j}^{k}>0}\right)u_{i,j}^{k+1}-\alpha u_{i,j-1}^{k+1}=\frac{u_{i+1,j}}{\Delta t}+Kg(e^{y_{j}})1_{g(e^{y_{j}})-u_{i,j}^{k}>0}.$ (E.5)
@@
[…]